OpenStudy (anonymous):

The path of a football kicked by a field goal kicker can be modeled by the equation y = –0.03x2 + 1.53x, where x is the horizontal distance in yards and y is the corresponding height in yards. Q1: What is the football’s maximum height? Round to the nearest tenth. Q2: How far is the football kicked?

OpenStudy (anonymous):

If you have a quadratic function in the form \(y=ax^2+bx+c\) the maximum is at \(x=-\frac{b}{2a}\). Can you tell what your a and b are?

OpenStudy (anonymous):

i believe a= -.03 and b=1.53 . but for some reason every time i answer the question it marks it wrong

OpenStudy (anonymous):

yes those are right, but -b/(2a) isn't the height. You have to plug what you got for that into the original formula and the y-value you get is the max height

OpenStudy (anonymous):

so you should get \(x=-\frac{1.53}{2(-0.03)}=25.5\). Then plug in 25.5 for x in the original

OpenStudy (anonymous):

oh ok the answer is 19.5 c:

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

For the second question, set y = 0 and solve for x. This one factors \(–0.03x^2 + 1.53x=0\)